A 65 mm diameter solid shaft is to be welded to a flat plate and is required to carry a torque of 1750 N-m. If fillet joint is used for welding what will be the minimum size of the weld when working shear stress is 66 MPa.


1
Expert's answer
2021-03-18T03:36:31-0400

If, h= Size of fillet weld

t= throat thickness

T= Torque applied

The maximum shear stress developed in the weld material is given by,

"\tau_s =\frac{2×T}{\pi ×t×d^2}"



Given as

"\tau_s=66 N/mm^2\\
T=1750 ×1000Nmm\\
d=65mm\\"

Solving for t, we get

"t=\frac{2T}{\pi×\tau_s×d^2}=\frac{2×1750×1000}{3.14×66×65^2}" =3.995=4 mm

Size of weld(h)="\frac{t}{0.707}= 5.66mm"